# UAV-Enabled ISAC Systems with Fluid Antennas

### v1

@@ IV-E Overall Algorithm
The AO-based algorithm for solving problem (P1) is outlined in Algorithm 1, which iteratively alternates among the optimization of the UAV trajectory, the beamforming design, the position of the transmit array, and the position of the receive array until the objective function value converges.
@@
IV-E1 Convergence Analysis
@@ V Numerical Results
This section presents numerical results to demonstrate the efficiency of FAs in the UAV-enabled ISAC system. As shown in Fig. 4, we consider an $800~\rm{m}\times 800~\rm{m}$ square area containing $M=6$ users and a single-point target, all spatially distributed within the region. Unless otherwise specified, the remaining system parameters are summarized in Table I [39, 30, 40].
@@
Channel power gain at reference distance $1\mathrm{m}$, $h_{0}$ -60 dB
Noise power at user receivers, $\sigma_{m}$ -90 dBm
Noise power at UAV receiver, $\sigma_{r}$ -90 dBm
RCS of target, $\varsigma_{\rm{T}}$ 1 $\rm{m}^{2}$
Length of transmission frame, $\tilde{L}$ 200
@@
Carrier wavelength, $\lambda$ 0.0107 $\mathrm{m}$
Minimum inter-antenna distance, $d_{\rm{min}}$ $0.5\lambda$
Maximum UAV speed, $V_{\rm{max}}$ 20m/s20\ $\mathrm${m/s}}
@@
Figure 4: UAV trajectory under different weight factors $\xi_{c}$.
@@
Figure 5: Sum rate of all users versus time slot index under different weight factors $\xi_{c}$.
@@
Figure 6: Target CRB versus time slot index under different weight factors $\xi_{c}$.
V-A Performance Analysis of ISAC System
@@
First, to comprehensively analyze the trade-off between communication and sensing performance, the system performance is evaluated under different weight factors $\xi_{c}$. Fig. 3 presents the convergence behavior of the proposed algorithm under various values of $\xi_{c}$. It is observed that the objective function, defined as the weighted sum of the communication rate and the inverse of the sensing CRB, converges within approximately 15 iterations. When $\xi_{c}=1$, the problem reduces to a communication-only case, achieving the highest objective function value. In contrast, when $\xi_{c}=0$, the optimization becomes a sensing-only case, resulting in the lowest objective function value. For intermediate values of $\xi_{c}$, the proposed algorithm provides a flexible balance between communication and sensing performance.
@@
As shown in Fig. 4, the optimized UAV trajectories vary considerably with different weight factors $\xi_{c}$. When $\xi_{c}=0$, the optimization mainly minimizes the sensing CRB, and the UAV tend to approach the target to enhance sensing performance. As $\xi_{c}$ increases, the objective gradually shifts towards improving the communication rate, and the UAV trajectories move closer to the user locations. Finally, when $\xi_{c}=1$, the UAV trajectories are primarily distributed around the user region to maximize the communication performance.
@@
Figs. 5 and 6 illustrate the variations of the total communication rate of all users and the target sensing CRB across the entire flight duration under different weight factors $\xi_{c}$. As shown in Fig. 5, increasing $\xi_{c}$ notably enhances the communication rate in most time slots, with the highest performance achieved at $\xi_{c}=0.75$. This indicates that the optimization tends to adjust the UAV trajectory toward regions favorable for user communications. However, this improvement comes at the expense of degraded sensing performance, as shown in Fig. 6, where the CRB values corresponding to $\xi_{c}=0.75$ are generally higher than those under other weight settings. When $\xi_{c}=0.25$, the system places greater emphasis on sensing accuracy, achieving substantially lower CRB values, but with a reduced communication rate. A moderate weight factor, such as $\xi_{c}=0.5$, provides a balanced trade-off between communication and sensing performance. These results demonstrate that adjusting $\xi_{c}$ enables a smooth trade-off between communication rate and sensing accuracy.
@@
Moreover, when $\xi_{c}=0.25$, the UAV trajectory shifts closer to the target during time slots 8-12 (as shown in Fig. 4), resulting in the lowest CRB values, which indicates that sensing accuracy is prioritized under this setting. However, during time slots 0-8 and 12-20, the UAV needs to fly toward or away from the target, leading to increased CRB values in these intervals.
@@
As shown in Fig. 7(a), in the sensing-only scenario, the UAV trajectory is biased toward the target location to maximize the reflected signal strength, resulting in a concentrated main lobe around the target and superior sensing gain. However, user coverage is neglected, leading to degraded communication performance when the UAV is far from the users. In contrast, in the communication-only scenario shown in Fig. 7(c), the trajectory is optimized closer to the user locations, resulting in stronger beam energy directed toward the users and improved link quality, but with significantly reduced sensing gain as the main lobe deviates from the target. Notably, the proposed ISAC scenario in Fig. 7(b) achieves a balanced performance: the UAV trajectory is distributed between the users and the target, while the beampattern provides considerable gain toward both the users and the target. This demonstrates that the proposed scheme effectively coordinates communication and sensing functionalities by optimizing the weighted sum of the communication rate and the sensing CRB.
@@
(b) Proposed ISAC scenario.
(c) Communication-only scenario.
Figure 7: Achieved beampattern gains in time slot 9 under three different scenarios.
@@
Figure 9: The objective function value versus maximum transmit power.
Figure 10: The objective function value versus number of transmit/receive antennas.
Figure 11: The objective function value versus flight altitude of the UAV.
@@ V-B Performance Analysis of FA System
  • PSO-based Scheme: The positions of both the transmit and receive FA arrays are optimized using the particle swarm optimization (PSO) algorithm, while all other variables are optimized using the proposed method [41].
@@
  • RPA Scheme: The positions of the transmit and receive FA arrays are generated randomly, with the remaining variables optimized using the proposed method.
@@
  • FPA Scheme: The positions of the transmit and receive FA arrays are configured as ULAs, while all other variables are optimized using the proposed method [30].
@@
Fig. 8 compares the convergence behavior of the proposed scheme with the three benchmark algorithms. It is evident that the proposed scheme exhibits a rapid growth of the objective function value in the early iterations and converges to the highest steady-state value after approximately 12 iterations, significantly outperforming all benchmarks. This improvement is attributed to the ability of the proposed method to fully exploit the reconfigurable properties of FA during optimization. The PSO-based scheme, where the transmit and receive positions are optimized using PSO, achieves better performance than RPA and FPA owing to its heuristic search capability. However, its final objective function value remains below that of the proposed method.
@@
Fig. 9 illustrates the objective function values achieved by different schemes under varying maximum transmit power $P_{\rm{max}}$. As observed, the objective function value consistently increases with $P_{\rm{max}}$ across all schemes, since higher transmit power enhances the received signal quality and thereby improves both communication and sensing performance. Nevertheless, the performance gaps among the schemes remain evident across the entire power range. The proposed scheme consistently achieves the highest objective function value under all power levels and maintains a significant advantage in the high-power regime.
@@
Fig. 10 presents the objective function values of different schemes versus the number of transmit/receive antennas. It can be observed that the objective function value increases monotonically with the number of antennas for all schemes. This is because a larger array provides higher beamforming gain and greater spatial diversity, which jointly enhance the communication rate and reduce the sensing accuracy. Nevertheless, the performance gap among the schemes remains consistent across the entire antenna range, with the proposed scheme achieving the highest objective function value under all considered antenna configurations.
@@
Fig. 11 illustrates the impact of UAV flight altitude on the objective function values for different schemes. It is observed that for all schemes, as the altitude increases from a low level, the objective function value first rises and then decreases, reaching its peak around 200-250 m. This trend can be explained as follows: a moderate increase in altitude enlarges the elevation angle for both communication and sensing links, thereby improving beam directivity, which benefits both throughput and sensing accuracy. However, when the altitude becomes excessively high, the increased distance leads to higher path loss, which dominates and results in overall performance degradation.
@@ VI Conclusion
In this paper, we propose a UAV-enabled ISAC system, in which a single UAV is equipped with transmit and receive FA arrays. By exploiting the mobility of antenna elements, the system provides additional spatial DoFs to enhance both communication and sensing performance. Specifically, a multi-objective optimization problem is formulated to maximize a weighted trade-off between communication and sensing performance through the joint optimization of the UAV trajectory, beamforming design, and the positions of the transmit and receive FA arrays. To solve this non-convex problem, an AO-based algorithm is developed to obtain a high-quality solution. Simulation results demonstrate that integrating FA technology into UAV-enabled ISAC systems achieves significantly improved performance compared to conventional FPAs. Future work will focus on extending the framework to multi-UAV cooperative ISAC networks with FA arrays, incorporating practical FA hardware constraints, and exploring learning-based optimization methods to further enhance real-time adaptability and efficiency.
@@ Appendix A Detailed Derivation Procedure for (14)
By substituting (88) and (92) into (13), (14) is finally obtained.
@@ Appendix B Detailed Derivation Procedure for (22)
By substituting (95) and (96) into (93), (22) is finally obtained.
@@ Appendix C Proof of the Rank-One Property of $\{\mathbf{W}_{m}^{\rm{opt}}[n]\}$
It is evident that problem (P2-2) constitutes a convex optimization problem satisfying Slater’s condition. Consequently, a zero duality gap exists between the primal and dual problems. Retaining only relevant terms, the Lagrangian of (P2-2) with respect to transmit beamforming matrices $\{\mathbf{W}{m}[n]\}$ is
@@ Appendix D Proof of Negative Semi-Definiteness of $\mathbf{S}^{\rm{I}}_{m}[n]$
$\small\varkappa_{m}[n]=\mathsf{diag}\left([\sqrt{|[\mathbf{w}_{m}[n]]_{1}|},\dots,\sqrt{|[\mathbf{w}_{m}[n]]_{N_{t}}|}]^{\mathsf{T}}\right)\in\mathbb{R}^{N_{t}\times N_{t}}.$ (101)
@@ Appendix E Proof of the Optimal solution for receiving FA position
Next, we determine the optimal index $j$, which specifies the position in the sequence where the placement shifts from the lower-bound cluster to the upper-bound cluster. If $j$ is too small, the transition occurs prematurely, reducing dispersion in the upper-bound region. Conversely, if $j$ is too large, the transition is delayed, limiting dispersion in the lower-bound region. Accordingly, the optimal index is determined as

### v2

@@ IV-E Overall Algorithm
The AO-based algorithm for solving problem (P1) is outlined in Algorithm 1, which iteratively alternates among the optimization of the UAV trajectory, the beamforming design, the transmit FA positions, an the receive FA positions until the objective function value converges.
@@
IV-E1 Initialization Analysis
For algorithm initialization, the UAV trajectory is set as a straight-line trajectory between the initial and final locations. The FA elements are uniformly distributed within the feasible region $0,D_{\mathrm{FA}}]$ with spacing $\frac{D_{\mathrm{FA}}}{K-1}$. The transmit beamforming vectors are initialized using the maximum ratio transmission (MRT) scheme [39].
IV-E2 Convergence Analysis
@@ V Numerical Results
This section presents numerical results to demonstrate the efficiency of FAs in the UAV-enabled ISAC system. As shown in Fig. 4, we consider an $800~\rm{m}\times 800~\rm{m}$ square area containing $M=6$ users and a single-point target, all spatially distributed within the region. Unless otherwise specified, the remaining system parameters are summarized in Table I.
@@
Channel power gain at reference distance $1\mathrm{m}$, $h_{0}$ -30 dB
Noise power at user receivers, $\sigma_{m}^{2}$ -90 dBm
Noise power at UAV receiver, $\sigma_{r}^{2}$ -90 dBm
RCS of target [40], $\varsigma_{\rm{T}}$ $10^{-6}$
Length of transmission frame [40], $\tilde{L}$ 256
@@
Carrier wavelength [41], $\lambda$ 0.0107 $\mathrm{m}$
Minimum inter-antenna distance [42], $d_{\rm{min}}$ $0.5\lambda$
Maximum UAV speed [30], $V_{\rm{max}}$ 20m/s20\ $\mathrm${m/s}}
@@
Figure 4: The UAV trajectory under different weight factors $\xi_{c}$.
@@
Figure 5: The sum rate of all users versus time slot index under different weight factors $\xi_{c}$.
@@
Figure 6: The target CRB versus time slot index under different weight factors $\xi_{c}$.
V-A Performance Analysis of UAV-enabled ISAC System
@@
First, to comprehensively analyze the trade-off between communication and sensing performance, the system performance is evaluated under different weight factors $\xi_{c}$. Fig. 3 presents the convergence behavior of the proposed algorithm under various values of $\xi_{c}$. It is observed that the objective function, defined as the weighted sum of the communication rate and the inverse of the sensing CRB, converges within approximately 15 iterations. When $\xi_{c}=1$, the problem reduces to a communication-only case, achieving the highest objective function value. In contrast, when $\xi_{c}=0$, the optimization becomes a sensing-only case, resulting in the lowest objective function value. For intermediate values of $\xi_{c}$, the proposed algorithm provides a flexible balance between the communication and sensing performance.
@@
As shown in Fig. 4, the optimized UAV trajectories vary considerably with different weight factors $\xi_{c}$. When $\xi_{c}=0$, the optimization mainly minimizes the sensing CRB, and the UAV tends to approach the target to enhance sensing performance. As $\xi_{c}$ increases, the objective gradually shifts towards improving the communication rate, and the UAV trajectories move closer to the user locations. Finally, when $\xi_{c}=1$, the UAV trajectories are primarily distributed around the user regions to maximize the communication performance.
@@
Figs. 5 and 6 illustrate the variations of the sum rate of all users and the target sensing CRB across the entire flight duration under different weight factors $\xi_{c}$. As shown in Fig. 5, increasing $\xi_{c}$ notably enhances the communication rate in each time slot, with the highest performance achieved at $\xi_{c}=0.6$. This indicates that the optimization tends to adjust the UAV trajectory toward regions favorable for user communications. However, this improvement comes at the expense of degraded sensing performance, as shown in Fig. 6, where the CRB values corresponding to $\xi_{c}=0.6$ are generally higher than those under other weight settings. For instance, when $\xi_{c}=0.2$, the system places greater emphasis on sensing accuracy, achieving substantially lower CRB values, but reduced communication rate. A moderate weight factor, such as $\xi_{c}=0.4$, provides a balanced trade-off between the communication and sensing performance. These results demonstrate that adjusting $\xi_{c}$ enables a better trade-off between communication rate and sensing accuracy.
@@
Moreover, when $\xi_{c}=0.2$, the UAV trajectory shifts closer to the target during time slots 7-14 (as shown in Fig. 4), resulting in the lowest CRB values, which indicates that sensing accuracy is prioritized under this setting. However, during time slots 0-7 and 14-20, the UAV needs to fly toward or away from the target, leading to increased CRB values in these intervals.
@@
As shown in Fig. 7(a), in the sensing-only scenario, the UAV trajectory is biased toward the target location to maximize the reflected signal strength, resulting in a concentrated main lobe around the target and superior sensing gain. However, user coverage is neglected, leading to degraded communication performance when the UAV is far from the users. In contrast, in the communication-only scenario shown in Fig. 7(b), the trajectory is optimized closer to the user locations, resulting in stronger beam energy directed toward the users and improved communication quality, but with significantly reduced sensing gain as the main lobe deviates from the target. Notably, the proposed ISAC scenario in Fig. 7(c) achieves a balanced performance: the UAV trajectory is distributed between the users and the target, while the beampattern provides considerable gain toward both the users and the target. This demonstrates that the proposed scheme effectively coordinates communication and sensing functionalities by optimizing the weighted sum of the communication rate and the sensing CRB.
@@
(b) Communication-only scenario.
(c) Proposed ISAC scenario.
Figure 7: The achieved beampattern gains in time slot 15 under three different scenarios.
@@
Figure 9: The objective function value versus the maximum transmit power.
Figure 10: The objective function value versus the number of transmit/receive antennas.
Figure 11: The objective function value versus the flight altitude of the UAV.
@@ V-B Performance Analysis of FA System
  • PSO-based Scheme: The transmit and receive FA positions are optimized using the particle swarm optimization (PSO) algorithm, while all other variables are optimized using the proposed method [43].
@@
  • RPA Scheme: The transmit and receive FA positions are generated randomly, with the remaining variables optimized using the proposed method.
@@
  • FPA Scheme: The transmit and receive FA positions are configured as ULAs, while all other variables are optimized using the proposed method [30].
@@
Fig. 8 compares the convergence behavior of the proposed scheme with the three benchmark algorithms. It is evident that the proposed scheme exhibits a rapid growth of the objective function value in the early iterations and converges to the highest steady-state value after approximately 12 iterations, significantly outperforming all benchmarks. This improvement is attributed to the ability of the proposed method to fully exploit the reconfigurable properties of FA during optimization. The PSO-based scheme, where the transmit and receive FA positions are optimized using PSO, achieves better performance than RPA and FPA owing to its heuristic search capability. However, its final objective function value remains below that of the proposed method.
@@
Fig. 9 illustrates the objective function values achieved by different schemes under varying maximum transmit power $P_{\rm{max}}$. As observed, the objective function values consistently increase with $P_{\rm{max}}$ across all schemes, since higher transmit power enhances the received signal quality and thereby improves both communication and sensing performance. Nevertheless, the performance gaps among the schemes remain evident across the entire power range. The proposed scheme consistently achieves the highest objective function value under all power levels and maintains a significant advantage in the high-power regime.
@@
Fig. 10 presents the objective function values of different schemes versus the number of transmit/receive antennas. It can be observed that the objective function values increase monotonically with the number of antennas for all schemes. This is because a larger array provides higher beamforming gain and greater spatial diversity, which jointly enhance the communication rate and reduce the sensing accuracy. Nevertheless, the performance gap among the schemes remains consistent across the entire antenna range, with the proposed scheme achieving the highest objective function value under all considered antenna configurations.
@@
Fig. 11 illustrates the impact of UAV flight altitude on the objective function values for different schemes. It is observed that for all schemes, as the altitude increases, the objective function value first rises and then decreases, reaching its peak around 200-250 m. This trend can be explained as follows: a moderate increase in altitude enlarges the elevation angle for both communication and sensing links, thereby improving beam directivity, which benefits both throughput and sensing accuracy. However, when the altitude becomes excessively high, the increased distance leads to higher path loss, which dominates and results in overall performance degradation.
@@ VI Conclusion
In this paper, we propose a UAV-enabled ISAC system, in which a single UAV is equipped with transmit and receive FA arrays. By exploiting the mobility of antenna elements, the system provides additional spatial DoFs to enhance both communication and sensing performance. Specifically, a multi-objective optimization problem is formulated to maximize a weighted trade-off between the communication and sensing performance through the joint optimization of the UAV trajectory, beamforming design, and the transmit and receive FA positions. To solve this non-convex problem, an AO-based algorithm is developed to obtain a high-quality solution. Simulation results demonstrate that integrating FA technology into UAV-enabled ISAC systems achieves significantly improved performance compared to conventional FPAs. Future work will focus on extending the framework to multi-UAV cooperative ISAC networks with FAs, incorporating practical FA hardware constraints, and exploring learning-based optimization methods to further enhance real-time adaptability and efficiency.
@@ Appendix A Detailed Derivation Procedure for (14)
By substituting (86) and (90) into (13), we can obtain (14).
@@ Appendix B Detailed Derivation Procedure for (22)
By substituting (93) and (94) into (91), we can obtain (22).
@@ Appendix C Proof of the Rank-One Property of $\{\mathbf{W}_{m}^{\rm{opt}}[n]\}$
It is evident that problem (P2-2) constitutes a convex optimization problem satisfying the Slater’s condition. Consequently, a zero duality gap exists between the primal and dual problems. Retaining only relevant terms, the Lagrangian of (P2-2) with respect to transmit beamforming matrices $\{\mathbf{W}_{m}[n]\}$ is
@@ Appendix D Proof of Negative Semi-Definiteness of $\mathbf{S}^{\rm{I}}_{m}[n]$
$\small\varkappa_{m}[n]=\mathsf{diag}\left(\sqrt{|[\mathbf{w}_{m}[n]]_{1}|},\dots,\sqrt{|[\mathbf{w}_{m}[n]]_{N_{t}}|}^{\mathsf{T}}\right)\in\mathbb{R}^{N_{t}\times N_{t}}.$ (99)
@@ Appendix E Proof of the Optimal solution for receiving FA position
Next, we determine the optimal index $j$, which specifies the position in the sequence where the placement shifts from the lower-bound cluster to the upper-bound cluster. If $j$ is too small, the transition occurs prematurely, reducing dispersion in the upper-bound region. Conversely, if $j$ is too large, the transition is delayed, limiting dispersion in the lower-bound region. Accordingly, the optimal index can be expressed as